On January 2 of the current year, Lem Corp. bought machinery under a contract that required a down payment of $10,000, plus 24 monthly payments of $5,000 each, for total cash payments of $130,000. The cash equivalent price of the machinery was $110,000. The machinery has an estimated useful life of ten years and estimated salvage value of $5,000. Lem uses straight-line depreciation. In its year-end income statement, what amount should Lem report as depreciation for this machinery?


The answer is $10,500. I get that. What would the journal entry look like for something like this?


1
Expert's answer
2021-03-17T18:32:00-0400

Solution:

The fair market value of $110,000 should be used to calculate depreciation. We cannot use the total cash payments of $130,000 since it includes interest as well as the principal.

The journal entry for machinery purchased will be as follows:

Particulars                                                          DR                          CR

Fixed assets                                                      110,000

        Cash                                                                                           10,000

        Account payable                                                                      100,000

(To record the purchase of machinery partly in cash and on credit)

The fair market value of $110,000 is used as the cost price of the machinery


Depreciation calculation:

Annual Depreciation expense = "\frac{Cost\; of\; the\; asset - Salvage\; value}{Useful\; life}"


= "\frac{110,000 - 5,000}{10} = \frac{105,000}{10} = 10,500"


Annual depreciation expense = $10,500


The journal entry will be as follows:

Particulars                                                          DR                          CR

Depreciation expense                                      10,500

       Accumulated Depreciation                                                      10,500

(To record the annual depreciation expense for the machinery)